# Detecting eclipsing double white dwarfs with electromagnetic and gravitational waves

### v1

@@ 2.3 Simulated DWDs
As mentioned above, we generate $N_{\rm fixed}$ DWDs as a fixed subset using different star formation histories for Model 1 and Model 2. To ensure adequate sampling of the short period end of DWDs, with periods of approximately several minutes, a relatively large $N_{\rm fixed}$ is necessary. After several trial experiments, we decided to set $N_{\rm fixed}=1.7\times 10^{8}$. We then scaled up the simulated fixed population of DWDs to represent the entire population expected in our Galaxy.
In addition to the intrinsic physical parameters discussed in Sec. 2.1, we need to assign values for the extrinsic geometric parameters to each pair of DWDs. Specifically, $R$ and $z$ are drawn according to Eq. 2 and Eq. 3 for the bulge and disk, respectively. The azimuth angle $\phi$ is drawn from the uniform distribution between $[0,2\pi]$. We transform $(R,z,\phi)$ to the Ecliptic, Galactocentric, ICRS coordinates to get ecliptic longitude, ecliptic latitude, distance (lon, lat, $d$), Galactic longitude, Galactic latitude, distance($l$, $b$, $d$), Right ascension, Declination, distance (RA, Dec, $d$) for the convenience of the next steps. Here $d$ represent the distance between the DWD and the observer.
The cosine of the orbital inclination $\iota$ follows a uniform distribution between $[0,1]$. Since the detection frequency of DWDs relevant to space-borne GW detectors is mainly in the millihertz range, we select DWDs with orbital periods of less than 30 hours for further analysis. In the end, we obtained final populations of Galactic DWDs with orbital periods less than 30 hours in Model 1 and Model 2, which amount to $6.8\times 10^{7}$ and $4.454\times 10^{7}$, respectively. In Figure 1, we show the distribution of a small representative subset of the simulated DWDs (grey squares) in our Galaxy, along with the ones detectable by EM and GW observations (see details in Section 3 and 4.).
Figure 1: Distribution of DWDs in the $x-z$ plane of the Galactic Cartesian coordinate system. Grey squares represent a small representative subset of the simulated Galactic DWDs. Blue squares represent DWDs detectable by TianQin ($\mathrm{SNR}>7$), and cyan squares represent DWDs detectable by Gaia and LSST.
@@ 3 EM observation
In this section, we utilize PHOEBE (the PHysics Of Eclipsing BinariEs; Prša, 2018; Prša & Zwitter, 2005; Prša et al., 2016; Horvat et al., 2018; Jones et al., 2020; Conroy et al., 2020) to generate light curves for the simulated DWDs in Section 2 and investigate the detection capabilities of DWDs based on the photometric performance of Gaia and LSST.
@@
The Gaia mission was launched in December 2013 and has since been in operation for over 10 years (Gaia Collaboration et al., 2016a). To obtain the most comprehensive information about our Galaxy, Gaia scans the entire sky and measures astrometry, photometry, radial velocities, and many astrophysical characteristics for billions of stars. Gaia has released several datasets containing approximately 1.8 billion sources brighter than 21 so far  (Gaia Collaboration et al., 2016b, 2018, 2021, 2022). These progressive data releases have seen impressive growth in the identification of white dwarf candidates within the catalog, increasing from 73,221 to 359,000 (Jiménez-Esteban et al., 2018; Gentile Fusillo et al., 2019, 2021b).
@@
The Vera C. Rubin Observatory is expected to start the Legacy Survey of Space and Time (LSST) in 2025. LSST will produce an abundance of images and data products that will contribute to exploring our solar system, galaxy, and universe over ten years of mission. The telescope will observe the southern celestial hemisphere (about 20000 square degrees) in six broad-band $(ugrizy)$ filters covering the wavelength range of 320 to 1050 nm. As the principal component of LSST, the main wide-fast-deep (WFD) survey will pay about 825 visits to each 9.6 square degree field summed over all six bands (LSST Science Collaboration et al., 2017; Bianco et al., 2021; Schwamb et al., 2023). For this study, we have employed simulated 10-year survey pointing databases generated by the Operation Simulator (OpSim; Naghib et al., 2019; Delgado & Reuter, 2016; Delgado et al., 2014) on NOIRLab’s Astro Data Lab [^1] https://datalab.noirlab.edu/ to get the regions of the sky covered by LSST. The shaded area in Figure 2 represents the sky coverage of the LSST-$r$ band.
@@
Figure 2: Distribution of detectable eclipsing DWDs in the LSST $r$-band across the celestial sphere from our simulation. The sky is shown in an equal-area Mollweide projection in equatorial coordinates (R.A. is increasing from left to right). The gray area indicates the sky region covered by LSST $r$-band.
Table 3: The GW amplitude $\mathcal{A}$ and SNR of 16 optical candidates of eclipsing DWDs. TianQin and LISA employ a nominal mission duration of four and five years, respectively.
       Name        $\mathcal{A}$        $\mathrm{SNR\ (TQ)}$        $\mathrm{SNR\ (LISA)}$        $\mathrm{SNR\ (TQ+LISA)}$
[…]
       NLTT $\mathrm{NLTT\ 11748}$        $8.605459\times 10^{-23}$        0.011357        0.016871        0.020962
       CSS $\mathrm{CSS\ 41177}$        $5.467304\times 10^{-23}$        0.02776843        0.08581914        0.09105031
[…]
@@ 3.1 Light curve simulation
PHOEBE can produce and fit light curves and radial velocity profiles of eclipsing binaries. In light curve simulation, the input parameters for the binary include: $r_{1}$, $r_{2}$, $\mathrm{T_{1}}$, $\mathrm{T_{2}}$, $a$, $d$, $\mathrm{P}$, $\iota$, and $e$, where $r_{1}$ and $r_{2}$ are the radii and $\mathrm{T_{1}}$ and $\mathrm{T_{2}}$ are the effective temperatures of each white dwarf, $a$ is the orbital semi-major axis, $d$ represents the distance of the DWD from the observer, $\mathrm{P}$ is the orbital period, $\iota$ is the orbital inclination, and $e$ is the orbital eccentricity. In addition to the primary influence of these intrinsic parameters on the light curve, we also consider secondary effects, such as limb-darkening and mutual irradiation effects. The limb-darkening coefficient is automatically obtained by interpolating the applicable tables based on the atmospheric parameters for ease of calculation (Prša et al., 2016). For the mutual irradiation effects, we adopt the default irradiation method ‘horvat’ (Horvat et al., 2019). However, we have not considered the surface distortion due to tides and rotation (ellipsoidal variation) in this work. Finally, we convert the output flux measured in $\mathrm{W/m^{2}}$ to apparent magnitudes for various bands in Gaia, LSST, and include the Gaussian noise from observational uncertainties and the extinction effect due to interstellar dust in our synthetic light curves.
Figure 3: Simulated light curves of an eclipsing DWDs in the Gaia $G$-band (top panel), the LSST-$r$ and -$i$ bands (bottom panel). Dash-dotted lines represent the simulated noiseless light curves. The number of simulated Gaia and LSST photometric data points are taken from Table 1.
According to the definition of Casagrande & VandenBerg (2018), Gaia magnitude is defined as
$G_{\zeta}=-2.5\log\bar{I}_{\zeta}+ZP_{\zeta}\,,$ (4)
where $\bar{I}_{\zeta}$ is the weighted mean flux in a given band $\zeta$ $(G_{\mathrm{BP}},G_{\mathrm{RP}},G)$ expressed in photoelectrons per second. This conversion from energy flux to photon flux follows the procedure outlined in Gaia DR2 documentation (Chapter 5; van Leeuwen et al., 2018). $ZP_{\zeta}$ represents the phoometric zero point for a given band. We calculate the Gaia $G$-band magnitude in the VEGAMAG system referring to the latest $ZP_{VEG}$ value of 25.6874, as presented in Table 3 of Riello et al. (2021). We use the photometric standard error of a single visit:
$\sigma_{G}[\mathrm{mag}]=1.2\times 10^{-3}\left(0.04895z_{x}^{2}+1.8633z_{x}+0.0001985\right)^{1/2}\,,$ (5)
@@
According to LSST Science Collaboration et al. (2009), we calculate the LSST magnitude in the AB magnitude system, expressed as follows:
$m_{\zeta}\equiv-2.5\log_{10}\left(\frac{F_{\zeta}}{F_{AB}}\right)\,,$ (6)
where $F_{\zeta}$ corresponds to the flux in a given band ($\zeta=ugrizy$) in units of Jansky (1 Jansky = $10^{-26}\,\mathrm{W\ Hz^{-1}}$ and $F_{AB}=3631\ \mathrm{Jy}$). The photometric error for a point source in a single visit can be written as
@@
We adopt the distance-dependent extinction model defined in Breivik et al. (2018) for the DWDs distributed in the Galactic disk and bulge.
In the disk, the extinction in the $V$ band
$A_{V}(d)_{\mathrm{disk}}=A_{V}(\infty)\tanh\left(\frac{d\sin(b)}{z_{h}}\right)\,,$ (8)
where $d$ is the distance, $b$ is the galactic latitude and the scale height $z_{h}=0.3~{}\mathrm{kpc}$.
In the bulge,
$A_{V}(d)_{\mathrm{bulge}}=A_{V}(\infty){\rm Erf}\left(\frac{d\sin(b)}{R_{h}}\right)\,,$ (9)
where $R_{h}=0.5~{}\mathrm{kpc}$ and $\rm{Erf}(\cdot)$ is the error function. $A_{V}(\infty)$ is defined in Nelemans et al. (2004) as follows:
$A_{\mathrm{V}}(\infty)=\begin{cases}0.165\frac{\left[\tan\left(50^{\circ}\right)-\tan(b)\right]}{\sin(b)}&\text{ for }b<50^{\circ}\,,\\ 0&\text{ for }b\geqslant 50^{\circ}\,.\end{cases}$ (10)
We transform the $V$-band extinction to the Gaia $G$-band and the LSST $(ugrizy)$ bands according to
$A(\lambda)=A_{V}\left(a(\tilde{x})+b(\tilde{x})/R_{V}\right)\,,$ (11)
where $R_{V}=3.1$ for our Galaxy. The coefficients $a(\tilde{x})$ and $b(\tilde{x})$ are functions of wavelength ($\tilde{x}=1/\lambda$), and $A_{V}$ is the extinction in the $V$ band for disk or for bulge (Cardelli et al., 1989).
Using the above equations, we simulate the light curves of eclipsing DWDs observable by Gaia and LSST. In Figure 3, we show the simulated light curves for an eclipsing DWD using the following parameters: $r_{1}=0.02980\ \mathrm{R}_{\sun}$, $r_{2}=0.0275\ \mathrm{R}_{\sun}$, $\mathrm{T_{1}}=26300\ \mathrm{K}$, $\mathrm{T_{2}}=9200\ \mathrm{K}$, $a=0.1227\ \mathrm{R}_{\sun}$, $d=1.6\ \mathrm{kpc}$, $\mathrm{P}=0.1466\ \mathrm{h}$, and $\iota=82^{\circ}$. The primary eclipse is clearly visible in the simulated light curves.
@@ 3.2 Source selection
From all simulated DWDs, we select detectable eclipsing DWDs in the Gaia and LSST surveys using criteria that include the survey limiting magnitudes, the target coordinates, the chi-square $\chi^{2}$ tests of the light curves, and the inclination angles of the DWDs.
For Gaia, we select DWDs that are brighter than magnitude 21 over the entire sky. For LSST, we select DWDs based on the sky coverage and the limiting magnitude of each bandpass. The sky map of LSST is shown in Figure 2, and the limiting magnitudes for each bandpass are listed in Table 1. To ensure the selected binaries are eclipsing, we also apply the criterion of $\cos\iota\leq(r_{1}+r_{2})/a$, where $\iota$, $r_{i}$, $a$ are the orbital inclination, radii of WDs, and orbital separation, respectively.
To distinguish eclipsing DWDs from non-variable stars with photometric fluctuations, we used a method similar to that of Korol et al. (2019) to select DWDs with $\chi^{2}>3$. Here,
$\chi^{2}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{G_{i}-G_{\text{mean}}}{\sigma_{i}}\right)^{2}\,,$ (12)
where $N$ is the total number of visits per star given in Table 1, $G_{i}$ represents the measured photometric magnitude from each visit, $\sigma_{i}$ represents the photometric error as stated in Equations 5 and 7, and $G_{\text{mean}}$ represents the mean photometric magnitude for all visits in the same bandpass. Based on the above selection criteria, the numbers of detectable eclipsing DWDs in our simulations are counted and summarized in Table 4.
Table 4: The number of detectable eclipsing DWDs in our Galaxy by Gaia and LSST in our simulations. For LSST, we list the number of detection for each bandpass in the LSST filters $(ugrizy)$, as well as the total number obtained when combining all bandpasses (the second last column). We also derive the total number of detections when combining observations from Gaia and LSST (the last column).
Gaia LSST total
@@
Figure 5: Distributions of the apparent magnitudes of the eclipsing DWDs in Model 1. The light orange and blue histograms represent samples from the Gaia $G$-band and the LSST-$r$ band, respectively. The orange and blue vertical dash lines represent the limiting magnitudes of the Gaia $G$-band and the LSST-$r$ band, respectively.
Figure 6: The corner plot for several parameters of eclipsing DWDs in the Gaia and LSST-$r$ band samples. The diagonal histograms show the distributions of the Galactic latitude, Declination (Dec), apparent magnitude, distance, and extinction. The scatter plots show the correlations between each pair of parameters.
@@
(b) LSST
Figure 7: Sky distributions of the eclipsing DWDs (Model 1) detectable by Gaia (left panel) and LSST (right panel).

### v2

@@ 2.3 Simulated DWDs
As mentioned above, we generate $N_{\rm fixed}$ DWDs as a fixed subset using different star formation histories for Model 1 and Model 2. To ensure adequate sampling of the short period end of DWDs, with periods of approximately several minutes, a relatively large $N_{\rm fixed}$ is necessary. After several trial experiments, we decided to set $N_{\rm fixed}=1.7\times 10^{8}$, so that the values of $N_{\rm Galaxy}/N_{\rm fixed}$ are 0.87, 1.72, and 5.17 for the bulge, thin disk, and thick disk, respectively. We then scaled up (scaled down for the bulge) the simulated fixed population of DWDs to represent the entire population expected in our Galaxy.
In addition to the intrinsic physical parameters discussed in Sec. 2.1, we need to assign values for the extrinsic geometric parameters to each pair of DWDs. Specifically, the radial distance $R$ and the height $z$ in the cylindrical polar coordinates are drawn according to the density profiles for the bulge and disk, respectively. The azimuth angle $\phi$ is drawn from the uniform distribution between $[0,2\pi]$. We transform $(R,z,\phi)$ to the Ecliptic, Galactocentric, ICRS coordinates to get ecliptic longitude, ecliptic latitude, distance (lon, lat, $d$), Galactic longitude, Galactic latitude, distance($l$, $b$, $d$), Right ascension, Declination, distance (RA, Dec, $d$) for the convenience of the next steps. Here $d$ represents the distance between the DWD and the observer.
The cosine of the orbital inclination $\iota$ follows a uniform distribution between $[0,1]$. Since the detection frequency of DWDs relevant to space-borne GW detectors is mainly in the millihertz range, we select DWDs with orbital periods of less than 30 hours for further analysis. Note that this period is much longer than those detectable by space-borne GW detectors and can potentially be reduced to a lower value without changing the following results. However, as shown in Sec.  5, a fraction of the DWDs detectable by Gaia and VRO will have periods on the order of hours. In the end, we obtained final populations of Galactic DWDs with orbital periods less than 30 hours in Model 1 and Model 2, which amount to $6.8\times 10^{7}$ and $4.454\times 10^{7}$, respectively. In Figure 1, we show the distribution of a small representative subset of the simulated DWDs (grey squares) in our Galaxy, along with the ones detectable by EM and GW observations (see details in Section 3 and 4.)
Figure 1: Distribution of DWDs in the $x-z$ plane of the Galactic Cartesian coordinate system. Grey squares represent a small representative subset of the simulated Galactic DWDs. Blue squares represent DWDs detectable by TianQin ($\mathrm{SNR}>7$), and cyan squares represent DWDs detectable by Gaia and VRO.
@@ 3 EM observation
Table 2: Science requirements for Gaia and VRO. The referenced parameters are from Gaia Collaboration et al. (2016a) and Ivezić et al. (2019), respectively.
Gaia VRO
Survey Area whole sky 20000 ${\rm deg}^{2}$
Single Visit Limiting Magnitude 21 $u$: 23.9; $g$: 25; $r$: 24.7;
$i$: 24; $z$: 23.3: $y$: 22.1
Total visits per sky patch 70 $u$: 56; $g$: 80; $r$: 184;
$i$: 184; $z$: 160: $y$: 160
Wavelength coverage 330–1050 nm 320–1050 nm ($ugrizy$)
Nominal mission lifetime 5 yr 10 yr
@@
In this section, we utilize PHOEBE (the PHysics Of Eclipsing BinariEs; Prša, 2018; Prša & Zwitter, 2005; Prša et al., 2016; Horvat et al., 2018; Jones et al., 2020; Conroy et al., 2020) to generate light curves for the simulated DWDs in detached phase in Section 2 and investigate the detection capabilities of DWDs based on the photometric performance of Gaia and VRO. The parameters of Gaia and VRO used in this paper are summarized in Table 2.
@@
The Gaia mission was launched in December 2013 and has since been in operation for over 10 years (Gaia Collaboration et al., 2016a). To obtain the most comprehensive information about our Galaxy, Gaia scans the entire sky and measures astrometry, photometry, radial velocities, and many astrophysical characteristics for billions of stars. Gaia has released several datasets containing approximately 1.8 billion sources brighter than 21 so far  (Gaia Collaboration et al., 2016b, 2018, 2021, 2022). These progressive data releases have seen impressive growth in the identification of white dwarf candidates within the catalog, increasing from 73,221 to 359,000 (Jiménez-Esteban et al., 2018; Gentile Fusillo et al., 2019, 2021b). So far, most of Gaia’s discoveries have been made by high precision measurements of proper motion, parallax, color, etc., which complement the discoveries made by light curves from SDSS, ASAS-SN, ATLAS, ZTF, GOTO, etc. Recently, part of the Gaia light curves have been released in Gaia DR3 along with its first catalogue of eclipsing binary candidates (Mowlavi et al., 2023).
@@
The Vera C. Rubin Observatory (VRO), formally known as Large Synoptic Survey Telescope (LSST), is expected to start in 2025. VRO will produce an abundance of images and data products that will contribute to exploring our solar system, galaxy, and universe over the nominal 10 year survey plan. The telescope will observe the southern celestial hemisphere (about 20000 square degrees) in six broad-band $(ugrizy)$ filters covering the wavelength range of 320 to 1050 nm. As the principal component of VRO, the main wide-fast-deep (WFD) survey will pay about 825 visits to each 9.6 square degree field summed over all six bands (LSST Science Collaboration et al., 2017; Bianco et al., 2021; Schwamb et al., 2023). The number of visits per sky patch, listed in Table 2, serves as the baseline for our simulation. The actual surveys, however, may differ to some extent, particularly with the inclusion of deep drilling fields. For this study, we have employed simulated 10-year survey pointing databases generated by the Operation Simulator (OpSim; Naghib et al., 2019; Delgado & Reuter, 2016; Delgado et al., 2014) on NOIRLab’s Astro Data Lab [^1] https://datalab.noirlab.edu/ to get the regions of the sky covered by VRO. The shaded area in Figure 2 represents the sky coverage of the VRO-$r$ band.
@@
Figure 2: Distribution of detectable eclipsing DWDs in the VRO $r$-band across the celestial sphere from our simulation. The sky is shown in an equal-area Mollweide projection in equatorial coordinates (R.A. is increasing from left to right). The gray area indicates the sky region covered by VRO $r$-band.
@@ 3.1 Light curve simulation
PHOEBE can produce and fit light curves and radial velocity profiles of eclipsing binaries. In light curve simulation, the input parameters for the binary include: $r_{1}$, $r_{2}$, $\mathrm{T_{1}}$, $\mathrm{T_{2}}$, $a$, $d$, $\mathrm{P}$, $\iota$, and $e$, where $r_{1}$ and $r_{2}$ are the radii and $\mathrm{T_{1}}$ and $\mathrm{T_{2}}$ are the effective temperatures of each white dwarf, $a$ is the orbital semi-major axis, $d$ represents the distance of the DWD from the observer, $\mathrm{P}$ is the orbital period, $\iota$ is the orbital inclination, and $e$ is the orbital eccentricity. In addition to the primary influence of these intrinsic parameters on the light curve, we also consider secondary effects, such as limb-darkening and mutual irradiation effects. The limb-darkening coefficient is automatically obtained by interpolating the applicable tables based on the atmospheric parameters for ease of calculation (Prša et al., 2016). For the mutual irradiation effects, we adopt the default irradiation method ‘horvat’ (Horvat et al., 2019). However, we have not considered the surface distortion due to tides and rotation (ellipsoidal variation) in this work. For the system shown in Fig. 3, the differences in the mean magnitudes between the case with and without the distortion are less than 0.01 mag for Gaia and VRO. Thus, ignoring the distortion would be appropriate for estimating the number of detectable DWDs. In contrast, a more accurate modeling of the light curve would be needed to estimate the parameters, especially for the short period system. Finally, we convert the output flux measured in $\mathrm{W/m^{2}}$ to apparent magnitudes for various bands in Gaia, VRO, and include the Gaussian noise from observational uncertainties and the extinction effect due to interstellar dust in our synthetic light curves.
Figure 3: Simulated light curves of an eclipsing DWDs in the Gaia $G$-band (top panel), the VRO-$r$ and -$i$ bands (bottom panel). Dash-dotted lines represent the simulated noiseless light curves. The number of simulated Gaia and VRO photometric data points are taken from Table 2. Note that the light curves shown here are phase folded. See the last paragraph of this subsection for the parameters used in this example.
For Gaia, we calculate the magnitude according to the definition of Casagrande & VandenBerg (2018) and use the photometric standard error of a single visit:
$\sigma_{G}=1.2\times 10^{-3}\left(0.04895z_{x}^{2}+1.8633z_{x}+0.0001985\right)^{1/2}\,,$ (2)
@@
Similarly, we calculate the VRO magnitude in the AB magnitude system as used in LSST Science Collaboration et al. (2009). The photometric error for a point source in a single visit can be obtained using the following relation:
@@
We adopt the distance-dependent extinction model defined in Breivik et al. (2018) for the DWDs distributed in the Galactic disk and bulge. We first calculate the $V$-band extinction and then convert it to the Gaia $G$-band and the VRO $(ugrizy)$ bands using the relationship developed in Cardelli et al. (1989).
Using the above prescription, we simulate the light curves of eclipsing DWDs observable by Gaia and VRO. In Figure 3, we show the simulated light curves with phase folded for an eclipsing DWD using the following parameters: $r_{1}=0.02980\ \mathrm{R}_{\sun}$, $r_{2}=0.0275\ \mathrm{R}_{\sun}$, $\mathrm{T_{1}}=26300\ \mathrm{K}$, $\mathrm{T_{2}}=9200\ \mathrm{K}$, $a=0.1227\ \mathrm{R}_{\sun}$, $d=1.6\ \mathrm{kpc}$, $\mathrm{P}=0.1466\ \mathrm{h}$, and $\iota=82^{\circ}$. The primary eclipse is clearly visible in the simulated light curves.
@@ 3.2 Source selection
From all simulated DWDs, we select detectable eclipsing DWDs in the Gaia and VRO surveys using criteria that include (i) the survey limiting magnitudes, (ii) the target coordinates, (iii) the $\chi^{2}$ tests of the light curves, and (iv) the inclination angles of the DWDs. For Gaia and VRO, a detectable eclipsing DWD must meet all these four criteria simultaneously.
For Gaia, we select DWDs that are brighter than magnitude 21 over the entire sky. For VRO, we select DWDs based on the sky coverage and the limiting magnitude of each bandpass. The sky map of VRO is shown in Figure 2, and the limiting magnitudes for each bandpass are listed in Table 2. For Gaia and VRO, to ensure the selected binaries are eclipsing, we also apply the criterion of $\cos\iota\leq(r_{1}+r_{2})/a$, where $\iota$, $r_{i}$, $a$ are the orbital inclination, radii of WDs, and orbital separation, respectively.
To distinguish eclipsing DWDs from non-variable stars with photometric fluctuations, we used the aforementioned $\chi^{2}$ test similar to that of Korol et al. (2019) to select DWDs with $\chi^{2}>3$. Here,
$\chi^{2}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{M_{i}-M_{\text{mean}}}{\sigma_{i}}\right)^{2}\,,$ (4)
where $N$ is the total number of visits per star given in Table 2, $M_{i}$ represents the measured photometric magnitude from each visit, $\sigma_{i}$ represents the photometric error as stated in Equations 2 and 3 for Gaia and VRO, respectively, and $M_{\text{mean}}$ represents the mean photometric magnitude for all visits in the same bandpass. Based on the above selection criteria, the numbers of detectable eclipsing DWDs in our simulations are counted and summarized in Table 3.
Table 3: The number of detectable eclipsing DWDs in our Galaxy by Gaia and VRO in our simulations. For VRO, we list the number of detection for each bandpass in the VRO filters $(ugrizy)$, as well as the total number obtained when combining all bandpasses (the second last column). We also derive the total number of detections when combining observations from Gaia and VRO (the last column).
Gaia VRO total
@@
Figure 5: Distributions of the mean apparent magnitudes of the eclipsing DWDs in Model 1. The light orange and blue histograms represent samples from the Gaia $G$-band and the VRO-$r$ band, respectively. The orange and blue vertical dash lines represent the limiting magnitudes of the Gaia $G$-band and the VRO-$r$ band, respectively.
Figure 6: The corner plot for several parameters of eclipsing DWDs in the Gaia and VRO-$r$ band samples. The diagonal histograms show the distributions of the Galactic latitude, Declination (Dec), apparent magnitude, distance, and extinction. The scatter plots show the correlations between each pair of parameters.
@@
(b) VRO
Figure 7: Sky distributions of the eclipsing DWDs (Model 1) detectable by Gaia (left panel) and VRO (right panel).